# On the Present Status of Inflationary Cosmology

### v1

@@ Abstract
We give a brief review of the basic principles of inflationary theory and discuss the present status of the simplest inflationary models that can describe Planck/BICEP/Keck observational data by choice of a single model parameter. In particular, we discuss the Starobinsky model, Higgs inflation, and $\alpha$-attractors, including the recently developed $\alpha$-attractor models with $SL(2,\mathbb{Z})$ invariant potentials. We also describe inflationary models providing a good fit to the recent ACT data, as well as the polynomial chaotic inflation models with three parameters, which can account for any values of the three main CMB-related inflationary parameters $A_{s}$, $n_{s}$ and $r$.
@@ 1 Introduction
The inflationary theory was invented at the beginning of the 80s as an attempt to solve numerous problems of the standard Big Bang cosmology Starobinsky:1980te ; Guth:1980zm ; Linde:1981mu ; Linde:1983gd ; Linde:1986fd . During the last 40 years, this theory evolved from a collection of bold scientific proposals to the cosmological paradigm describing the origin of the universe and the formation of its large-scale structure.
The first realistic model of inflationary type was proposed by Starobinsky Starobinsky:1980te . However, it did not attempt to solve the homogeneity, isotropy, horizon, and flatness problems, which were the main goals of the inflationary scenario proposed by Guth Guth:1980zm . In the beginning, the Starobinsky model stayed slightly outside the mainstream of inflationary cosmology, but then it was realized that its predictions derived by Mukhanov and Chibisov Mukhanov:1981xt provided a perfect match to the WMAP and Planck data.
@@ 2 Chaotic inflation
Here $k=-1,0,1$ for an open, flat or closed universe respectively. We work in units $M_{p}^{-2}=8\pi G=1$.
@@ 3 Initial conditions for inflation
The main problem of the old Big Bang theory was that the cosmological evolution was supposed to be approximately adiabatic, and the total number of particles in the universe was supposed to be approximately conserved, changing by no more than one or two orders of magnitude. From this fact, one could show that at the Planckian time $t=O(1)$ when the density of the hot universe was Planckian $\rho=O(1)$, it consisted of about $10^{{90}}$ causally disconnected parts of a Planckian size $l=O(1)$, each of which contained only $O(1)$ elementary particles. In this situation, large-scale homogeneity and isotropy of the universe would appear miraculous, and the origin of the exponentially large number of particles of the universe would seem equally mysterious.
@@
[…]
This conclusion matches the results of the investigation of the quantum creation of the universe “from nothing”, which indicated that this process is possible at $V=O(1)$, but the probability becomes exponentially suppressed if $V\ll 1$ Linde:1983mx ; Linde:1983cm ; Linde:1984ir ; Vilenkin:1984wp . However, later it was realized that the quantum creation of a compact open or flat universe is not exponentially suppressed even if $V\ll 1$ Coule:1999wg ; Linde:2004nz . After a long stage of inflation, the universe becomes enormous and locally isotropic independently of its topology. Therefore, quantum creation of a compact open or flat universe is one of the simplest ways to account for inflation with inflationary potentials with $V\ll 1$. This comment also applies to the possibility of inflation with $\phi\ll 1$.
Yet another possibility to describe inflation at $V\ll 1$ and/or $\phi\ll 1$ is to consider multi-field inflationary models. For example, one may consider a model
@@
where $V(\phi)$ is a plateau potential of the type which appears in the $\alpha$-attractor models to be considered in this paper. Inflation driven by the field $\chi$ may begin at ${m^{2}\over 2}\chi^{2}=O(1)$, as in the simplest chaotic inflation model discussed above. If the field $\phi$ initially was at the plateau of its potential $V(\phi)$, it did not move until the end of the first stage of inflation driven by the field $\chi$. After that, the stage of inflation driven by the field $\phi$ begins, and if this stage is long enough, it determines the formation of the large-scale structure and CMB perturbations in the observable part of the universe. Various versions of this scenario have been discussed in Linde:1987yb ; Linde:2014nna ; Linde:2017pwt ; Corman:2022alv . We will return to it in this paper, in the context of the Starobinsky model.
@@ 4 Can we test inflation?
    The universe must be almost exactly flat, $\Omega=1$. Most inflationary models predict $\Omega=1\pm 10^{{-4}}$. 45 years ago, when we only knew that $\Omega=O(1)$. In the mid-90s, the consensus was that $\Omega=0.3$, until the discovery of dark energy, confirming inflation. The latest observational results show that $\Omega=1\pm 10^{{-2}}$, in agreement with inflationary predictions.
@@
    These perturbations are approximately Gaussian. In non-inflationary models, the parameter $f_{NL}$ describing the level of the so-called local non-Gaussianity can be as large as $O(10^{4})$, but it is predicted to be $O(1)$ in all single-field inflationary models. Prior to the Planck 2013 data release, there were rumors that $f_{NL}=O(30)$, which would rule out most single-field inflationary models. The latest results confirm that $|f_{NL}|\lesssim 5$.
@@
    Inflationary perturbations should have a nearly flat spectrum with the spectral index $n_{s}$ close to $1$. The latest results show that $n_{s}\sim 0.96-0.98$.
@@
    Inflationary perturbations should produce specific peaks in the spectrum of CMB radiation. Indeed, many such peaks have been found. The existence of such peaks ruled out alternative theories of the origin of the large-scale structure of the universe based on cosmic strings and textures.
@@
The next stage of investigation, discussed below, is to examine the predictions of specific inflationary models and compare them with each other.
Figure 1: ACT DR6 and Planck PR3 (Planck Collaboration 2020b) combined TT (top), EE (middle), and TE (bottom) power spectra. The gray lines show the joint ACT and Planck ${\rm\Lambda CDM}$ best-fit power spectra.
@@ 5 Constraints on inflationary models
The rapid progress in this direction can be illustrated by Fig. 2. It shows the broad range of allowed values of $n_{s}$ and $r$ in the 7-year WMAP data release, superimposed with the constraints in the Planck 2018 data release. As one can see, many inflationary models compatible with WMAP have been ruled out by Planck 2018.
@@
The P-ACT-LB results are illustrated by Fig. 10 in ACT:2025tim ; we reproduce it here in Fig. 4. The authors of Louis:2025tst conclude, in particular, that the P-ACT-LB constraints on $n_{s}$ disfavor the Starobinsky model Starobinsky:1980te at $\gtrsim 2\sigma$. This is a strong and rather unexpected result, very different from the results obtained before.
Note that the constraint $n_{s}=0.9743\pm 0.0034$ shown in Fig. 10 in ACT:2025tim was obtained after taking into account DESI Y1 results, but assuming $\Omega_{m}+\Omega_{\Lambda}=1$, i.e. in the $\Lambda$CDM context, see Table 5 in Louis:2025tst . On the other hand, more recent DESI Y2 results DESI:2025gwf strongly disfavor $\Lambda$CDM, which was used in the derivation of this constraint. The authors of DESI:2025gwf say: “For our baseline DESI DR2 BAO+Planck PR4+ACT likelihood combination, the preference for evolving dark energy over a cosmological constant is about $3\sigma$, increasing to over $4\sigma$ with the inclusion of Type Ia supernova data.”
Perhaps one should wait until the more recent DESI results are taken into account and also until the next SPT data release, which is expected soon. Previously, the SPT results favored much lower values of $n_{s}$ SPT-3G:2024atg .
Instead of making a judgement call in this respect, we will consider in this paper the simplest inflationary models consistent with the Planck/BICEP/Keck results shown in Fig. 3, and then we will discuss the models matching the results of Planck, ACT, and DESI (P-ACT-LB) shown in Fig. 4.
But before doing it, we will describe a basic chaotic inflation model with a polynomial potential with 3 parameters, which can describe any value of the amplitude of the spectrum of scalar perturbations $A_{s}$, of the spectral index $n_{s}$ and of the tensor-to-scalar ratio $r$. This may help us to look at the rest of the paper from a different perspective.
@@ 6 Polynomial chaotic inflation
According to BICEP:2021xfz , the simplest models of chaotic inflation $V(\phi)\sim\phi^{n}$ are ruled out by observations. However, it is not widely recognized that one can easily describe all presently available CMB-related data by models with simple polynomial potentials with only three parameters, see Fig. 2 in Destri:2007pv , Fig. 2 in Nakayama:2013jka , and Fig. 3 in Kallosh:2014xwa .
@@
The values of $n_{s}$ and $r$ for $N_{e}=55$ in this model are shown in Fig. 6, which is the lower part of Fig. 3 in Kallosh:2014xwa . Let us consider, for example, the yellow line with many small red dots. This line corresponds to $b=0.34$. The red dots, from the top down, correspond to $a$ changing from $a=0$ to $a=0.13$.
For completeness, we give here the results for $N_{e}=60$. For $b=0.34$, $a=0.13$ the predictions nearly coincide with the predictions of the Starobinsky model and the Higgs inflation model, as well as with the predictions of $\alpha$-attractors with $\alpha=1$:  $n_{s}=0.967$, $r\approx 3\times 10^{{-3}}$. For $a=0.14$, the value of $n_{s}$ becomes $0.973$, matching the ACT results shown in Fig. 4. A further increase of $a$ gradually increases the value of $n_{s}$, which reaches $n_{s}=1.006$ for $a=0.17$.
Figure 6: Predictions for $n_{s}$ and $r$ at $N_{e}=55$ in the model with $V(\phi)={m^{2}\phi^{2}\over 2}\,\bigl{(}1-a\phi+a^{2}b\,\phi^{2})\bigr{)}^{2}$ for $b=0.34$. The red dots, from the top down, correspond to $a=0$, 0.03, 0.05, 0.1, 0.13. For $a=0$, one recovers the predictions for the simplest chaotic inflation model with a quadratic potential; for $a=0.13-0.14$, the predictions almost exactly coincide with the predictions of the Starobinsky model and the Higgs inflation model.
@@
Thus we have a polynomial chaotic inflation model (8) with 3 parameters, $m$, $a$ and $b$, which can match any values of the 3 main inflationary parameters: the amplitude of scalar perturbations $A_{s}$, the spectral index $n_{s}$, and the tensor-to-scalar ratio $r$. The problem of initial conditions for inflation in this model is trivially solved in the same way as in the simplest models of chaotic inflation. This model does not require any profound modifications of the fundamental theory. There is no need to consider modifications of the general theory of relativity, as in $R+R^{2}$ Starobinsky model, which we will discuss shortly. There is no need to invoke non-minimal coupling of the inflaton field to gravity $\xi\phi^{2}R$ with $\xi\gg 1$, as in the Higgs inflation, or modify its kinetic terms, as in the theory of $\alpha$-attractors. This version of inflationary theory just works. It provides a basic framework that can describe any  inflationary parameters $A_{s}$, $n_{s}$, and $r$. In particular, a proper choice of parameters $a$ and $b$ allows one to match any values of $n_{s}$ and $r$ favored either by Planck/BICEP/Keck or by ACT.
Then why do we even need to continue discussing other models? The main reason is that some of them can match the Planck/BICEP/Keck data using just a single parameter required to fix the amplitude of the perturbations $A_{s}$. Moreover, some of these models (attractors) have universal predictions that are very stable with respect to the choice of the inflationary potential. Finally, some of the models may have better motivation from the point of view of quantum gravity, supergravity, and/or string theory. Such models, in addition to explaining the data, may provide us with important insights into fundamental physics.
@@ 7 Starobinsky model
According to the original version of the Starobinsky model, the universe at $t=0$ was in an unstable de Sitter state with the Hubble constant $H$ many orders of magnitude smaller than the Planck mass. In this sense, the original model was similar to new inflation, with inflation beginning at the top of the potential.
Since de Sitter space is non-singular, one should be able to continue this solution to $t\to-\infty$. However, an unstable de Sitter state could not survive from $t=-\infty$ to $t=0$. Another problem was to find a realistic model with proper parameters $k_{2}$ and $k_{3}$, which turned out to be very difficult.
The streamlined formulation of the Starobinsky model (9) was proposed in 1985 Kofman:1985aw , where Starobinsky explained why it was difficult to develop a realistic version of the original model Starobinsky:1980te . A transition from (9) to (10) took a few more years and the efforts of several different authors Whitt:1984pd ; Barrow:1988xh ; Maeda:1987xf ; Coule:1987wt .
@@ 8 Higgs inflation and $\xi$-attractors
In general, both of these models match the Planck data very well, and the fact that these two very different models have nearly identical predictions was enigmatic. This stimulated attempts to identify the reasons for such a coincidence and to look for other models that could have similar properties.
@@
The values of $n_{s}$ and $r$ as a function of $\xi$ are shown in Fig. 7 for $V(\phi)\sim\phi^{8},\phi^{6},\phi^{4},\phi^{2},\phi,\phi^{2/3}$. The dots in the right panel correspond to $\xi=1/e$, 1, and $e$, from the top down. As one can see, the convergence of the predictions for the potentials $V(\phi)\sim\phi^{8},\phi^{6},\phi^{4}$ to the attractor point $n_{s}=1-{2\over N_{e}}$, $r={12\over N^{2}_{e}}$ is very fast, the attractor point is reached already at $\xi=O(1)$. Meanwhile for the models $\phi^{2},\phi,\phi^{2/3}$ this limit is reached only for relatively large $\xi$. Thus, for these models it makes sense to consider a more natural possibility $\xi=O(1)$. This consideration will be important for us later, see section 13.1.

### v3

@@ Abstract
We give a brief review of the basic principles of inflationary theory and discuss the present status of the simplest inflationary models which can describe Planck/BICEP/Keck observational data by choice of a single model parameter. In particular, we discuss the Starobinsky model, Higgs inflation, and $\alpha$-attractors, including the recently developed $\alpha$-attractor models with $SL(2,\mathbb{Z})$ invariant potentials. We also describe inflationary models providing a good fit to the recent ACT data, as well as the polynomial chaotic inflation models with three parameters, which can account for any values of the three main CMB-related inflationary parameters $A_{s}$, $n_{s}$ and $r$.
@@ 1 Introduction
The inflationary theory was invented at the beginning of the 80s to address numerous problems of the standard Big Bang cosmology Starobinsky:1980te ; Guth:1980zm ; Linde:1981mu ; Linde:1983gd ; Linde:1986fd . Over the last 40 years, this theory has evolved from a collection of bold scientific proposals into the cosmological paradigm that describes the origin of the universe and the formation of its large-scale structure.
The first realistic model of inflationary type was proposed by Starobinsky Starobinsky:1980te . However, it did not attempt to solve the homogeneity, isotropy, horizon, and flatness problems, which were the main goals of the inflationary scenario proposed by Guth Guth:1980zm . Therefore, initially, the Starobinsky model remained slightly outside the mainstream of inflationary cosmology; however, it was later realized that the predictions of this model, derived by Mukhanov and Chibisov Mukhanov:1981xt , provided a perfect match to the WMAP and Planck data.
@@
In this paper, we will describe some of the basic and most popular inflationary models, discuss their cosmological predictions, and their status in view of the latest observational data.
@@ 2 Chaotic inflation
Here k=1,0,1k=-1,0,1 for an open, flat or closed universe respectively. We work in units $M_{p}^{-2}=8\pi G=1$.
@@ 3 Initial conditions for inflation
The primary issue with the old Big Bang theory was that cosmological evolution was supposed to be approximately adiabatic, and the total number of particles in the universe was supposed to be approximately conserved, changing by no more than one or two orders of magnitude. From this fact, one could show that at the Planckian time $t=O(1)$ when the density of the hot universe was Planckian $\rho=O(1)$, it consisted of about $10^{{90}}$ causally disconnected parts of a Planckian size $l=O(1)$, each of which contained only $O(1)$ elementary particles. In this situation, large-scale homogeneity and isotropy of the universe would appear miraculous, and the origin of the exponentially large number of particles of the universe would seem equally mysterious.
@@
During the last decade, the problem of initial conditions for inflation was thoroughly investigated by sophisticated numerical methods of the general theory of relativity Carrasco:2015rva ; East:2015ggf ; Kleban:2016sqm ; Clough:2016ymm ; Linde:2017pwt ; Clough:2017efm ; Aurrekoetxea:2019fhr ; Creminelli:2020zvc ; Joana:2020rxm ; Joana:2022pzo ; Corman:2022alv ; Elley:2024alx ; Aurrekoetxea:2024mdy ; Joana:2024ltg . All of these papers not only confirmed the validity of the basic principles of chaotic inflation, but also found that inflation can naturally begin even in the models with $V(\phi)\ll 1$. [^1] A notable exception is Ref. Garfinkle:2023vzf claiming that if the initial value of potential energy density of the inflaton ${m^{2}\over 2}\phi^{2}$ is many orders of magnitude smaller than $O(1)$ and its kinetic and gradient energy density and the Weyl tensor are $O(1)$, inflation does not last sufficiently long, $N_{e}<60$. It was argued in Garfinkle:2023vzf that all other authors Carrasco:2015rva ; East:2015ggf ; Kleban:2016sqm ; Clough:2016ymm ; Linde:2017pwt ; Clough:2017efm ; Aurrekoetxea:2019fhr ; Creminelli:2020zvc ; Joana:2020rxm ; Joana:2022pzo ; Corman:2022alv ; Elley:2024alx ; Aurrekoetxea:2024mdy ; Joana:2024ltg reached an opposite conclusion because they did not consider sufficiently general initial conditions. However, the real reason was much simpler: The initial value of $\phi$ considered in Garfinkle:2023vzf did not satisfy the basic requirement (6). In their next paper Ijjas:2024oqn , the authors increased $\phi_{0}$ and no longer claimed that $N_{e}<60$.
The simplest possibility to describe inflation at $V\ll 1$ and/or $\phi\ll 1$ is to consider multi-field inflationary models. For example, one may consider a model
@@
where $V(\phi)$ is a plateau potential of the type which appears in the $\alpha$-attractor models to be considered in this paper. Inflation driven by the field $\chi$ may begin at ${m^{2}\over 2}\chi^{2}=O(1)$, as in the simplest chaotic inflation model discussed above. If the field $\phi$ initially was at the plateau of its potential $V(\phi)$, it did not move until the end of the first stage of inflation driven by the field $\chi$. After that, the stage of inflation driven by the field $\phi$ begins, and if this stage is long enough, it determines the formation of the large-scale structure and CMB perturbations in the observable part of the universe. Various versions of this scenario have been discussed in Linde:1987yb ; Linde:2014nna ; Linde:2017pwt ; Corman:2022alv . We will revisit this topic in this paper, within the context of the Starobinsky model.
@@ 4 Can we test inflation?
    The universe must be almost exactly flat, $\Omega=1$. Most inflationary models predict $\Omega=1\pm 10^{{-4}}$. 45 years ago, when we only knew that $\Omega=O(1)$. In the mid-1990s, the consensus was that $\Omega=0.3$, until the discovery of dark energy, which allowed $\Omega=1$, consistent with inflation. The latest observational results show that $\Omega=1\pm 10^{{-2}}$, in agreement with inflationary predictions.
@@
    These perturbations are approximately Gaussian. In non-inflationary models, the parameter $f_{NL}$ describing the level of the so-called local non-Gaussianity can be as large as $O(10^{4})$, but it is predicted to be $O(1)$ in all single-field inflationary models. Prior to the Planck 2013 data release, rumors circulated that $f_{NL}=O(30)$, which would rule out most single-field inflationary models. The latest results confirm that $|f_{NL}|\lesssim 5$.
@@
    Inflationary perturbations should have a nearly flat spectrum with the spectral index $n_{s}$ close to 11. The latest results show that $n_{s}\sim 0.96-0.98$.
@@
    Inflationary perturbations should produce specific peaks in the spectrum of CMB radiation. Indeed, many such peaks have been found. The existence of such peaks ruled out alternative theories of the universe’s large-scale structure in the context of cosmic string and texture theories.
@@
The next stage of investigation, discussed below, involves examining the predictions of specific inflationary models and comparing them with one another.
Figure 1: ACT DR6 and Planck PR3 (Planck Collaboration 2020b) combined $T$ (top), $E$ (middle), and TE (bottom) power spectra. The gray lines show the joint ACT and Planck $\rm\Lambda CDM$} best-fit power spectra.
@@ 5 Constraints on inflationary models
The rapid progress in this direction can be illustrated by Fig. 2. It shows the broad range of allowed values of $n_{s}$ and $r$ in the 7-year WMAP data release, superimposed with the constraints in the Planck 2018 data release. As one can see, many inflationary models compatible with WMAP have been ruled out by the Planck 2018 results.
@@
The P-ACT-LB results are illustrated by Fig. 10 in ACT:2025tim ; we reproduce it here in Fig. 4. The authors of Louis:2025tst conclude, in particular, that the P-ACT-LB constraints on $n_{s}$ disfavor the Starobinsky model Starobinsky:1980te at $\gtrsim 2\sigma$. This is a strong and unexpected result, significantly different from the previous results. The latest SPT data, in combination with all other CMB results plus DESI DR2, give $n_{s}=0.9728\pm 0.0027$ SPT-3G:2025bzu , which is consistent with the ACT-DESI results.
[…]
Instead of making a judgment call in this respect, we will consider in this paper the simplest inflationary models consistent with the Planck/BICEP/Keck results shown in Fig. 3, and then we will discuss the models matching the results of Planck, ACT, and DESI (P-ACT-LB) shown in Fig. 4.
But before doing it, we will describe a basic chaotic inflation model with a polynomial potential with 3 parameters, which can describe any value of the amplitude of the spectrum of scalar perturbations $A_{s}$, of the spectral index $n_{s}$ and of the tensor-to-scalar ratio $r$. This may help us view the rest of the paper from a different perspective.
@@ 6 Polynomial chaotic inflation
According to BICEP:2021xfz , the simplest models of chaotic inflation $V(\phi)\sim\phi^{n}$ are ruled out by observations. However, it is not widely recognized that one can easily describe all presently available CMB-related data using models with simple polynomial potentials and only three parameters, as seen in Fig. 2 of Destri:2007pv , Fig. 2 of Nakayama:2013jka , and Fig. 3 of Kallosh:2014xwa .
@@
The values of $n_{s}$ and $r$ for $N_{e}=55$ in this model are shown in Fig. 6, which is the lower part of Fig. 3 in Kallosh:2014xwa . Let us consider, for example, the yellow line with many small red dots. This line corresponds to $b=0.34$. The red dots, from top to bottom, correspond to $a$ changing from $a=0$ to $a=0.13$.
For completeness, we present the results for $N_{e}=60$ here. For $b=0.34$, $a=0.13$, the predictions nearly coincide with the predictions of the Starobinsky model and the Higgs inflation model, as well as with the predictions of $\alpha$-attractors with $\alpha=1$: $n_{s}=0.967$, $r\approx 3\times 10^{{-3}}$. For $a=0.14$, the value of $n_{s}$ becomes $0.973$, matching the ACT results shown in Fig. 4. A further increase of $a$ gradually increases the value of $n_{s}$, which reaches $n_{s}=1.006$ for $a=0.17$.
Figure 6: Predictions for $n_{s}$ and $r$ at $N_{e}=55$ in the model with $V(\phi)={m^{2}\phi^{2}\over 2}\,\bigl(1-a\phi+a^{2}b\,\phi^{2})\bigr)^{2}$ for $b=0.34$. The red dots, from the top down, correspond to $a=0$, 0.03, 0.05, 0.1, 0.13. For $a=0$, one recovers the predictions for the simplest chaotic inflation model with a quadratic potential; for a=0.130.14a=0.13-0.14, the predictions almost exactly coincide with the predictions of the Starobinsky model and the Higgs inflation model.
@@
Thus, we have a polynomial chaotic inflation model (8) with 3 parameters, $m$, $a$ and $b$, which can match any values of the 3 main inflationary parameters: the amplitude of scalar perturbations $A_{s}$, the spectral index $n_{s}$, and the tensor-to-scalar ratio $r$. The problem of initial conditions for inflation in this model is trivially solved in the same way as in the simplest models of chaotic inflation. This model does not require any profound modifications of the fundamental theory. There is no need to consider modifications of the general theory of relativity, as in $R+R^{2}$ Starobinsky model, which we will discuss shortly. There is no need to invoke non-minimal coupling of the inflaton field to gravity $\xi\phi^{2}R$ with $\xi\gg 1$, as in the Higgs inflation, or modify its kinetic terms, as in the theory of $\alpha$-attractors. This version of inflationary theory just works. It provides a basic framework that can describe any  inflationary parameters $A_{s}$, $n_{s}$, and $r$. In particular, a proper choice of the parameters $a$ and $b$ allows one to match any values of $n_{s}$ and $r$ favored by either Planck/BICEP/Keck or ACT.
Then why do we even need to continue discussing other models? The main reason is that some of them can match the Planck/BICEP/Keck data using just a single parameter required to fix the amplitude of the perturbations $A_{s}$. Moreover, some of these models (attractors) have universal predictions that are very stable with respect to the choice of the inflationary potential. Finally, some of the models may have better motivation from the perspective of quantum gravity, supergravity, and/or string theory. Such models, in addition to explaining the data, may provide us with important insights into fundamental physics.
@@ 7 Starobinsky model
According to the original version of the Starobinsky model, the universe at $t=0$ was in an unstable de Sitter state with the Hubble constant $H$ many orders of magnitude smaller than the Planck mass. In this sense, the original model was similar to the new inflation, with inflation beginning at the top of the potential.
Since de Sitter space is non-singular, one should be able to continue this solution to $t\to$-$\infty$. However, an unstable de Sitter state could not survive from $t=$-$\infty$ to $t=0$. Another problem was finding a realistic model with proper parameter values $k_{2}$ and $k_{3}$, which proved very difficult.
The streamlined formulation of the Starobinsky model (9) was proposed in 1985 Kofman:1985aw , where Starobinsky explained why it was difficult to develop a realistic version of the original model Starobinsky:1980te . A transition from (9) to (10) took a few more years and the efforts of several authors Whitt:1984pd ; Barrow:1988xh ; Maeda:1987xf ; Coule:1987wt .
@@ 8 Higgs inflation and $\xi$-attractors
In general, both models match the Planck data very well, and the fact that these very different models have nearly identical predictions was enigmatic. This prompted attempts to identify the reasons for this coincidence and to seek other models with similar properties.
@@
Due to the non-minimal coupling of the scalar field to gravity, we will refer to this form of the theory as the Jordan frame. In order to transform to the canonical Einstein frame, we redefine the metric $g_{\mu\nu}\rightarrow\Omega(\phi)^{-1}g_{\mu\nu}$. This brings the Lagrangian to the Einstein-frame form:
${\mathcal{L}_{\rm E}\over\sqrt{-g}}=\tfrac{1}{2}R-\tfrac{1}{2}\Big(\Omega(\phi)^{-1}+\tfrac{3}{2}(\log\Omega(\phi))^{\prime 2}\Big)(\partial\phi)^{2}-V_{E}(\phi)\,,\quad{\rm where~~}V_{E}(\phi)=\frac{V_{J}(\phi)}{\Omega(\phi)^{2}}\,.$ (15)
The kinetic terms in (15) give rise to the following definition of the canonical scalar field $\varphi$:
$\frac{\partial\varphi}{\partial\phi}=\sqrt{\Omega(\phi)^{-1}+\tfrac{3}{2}(\log\Omega(\phi))^{\prime 2}}\,$, (16)
Many models discussed in this paper have non-minimal canonical terms. Sometimes the transition to the canonical variables $\varphi$ is simple, sometimes it is not. A general method to investigate such models and to find $n_{s}$ and $r$ beyond the slow-roll approximation can be found in Appendix A where we describe this method for convenience.
The values of $n_{s}$ and $r$ as a function of $\xi$ are shown in Fig. 7 for $V(\phi)\sim\phi^{8},\phi^{6},\phi^{4},\phi^{2},\phi,\phi^{2/3}$. The dots in the right panel correspond to $\xi=1/e$, 1, and $e$, from the top down. As one can see, the convergence of the predictions for the potentials $V(\phi)\sim\phi^{8},\phi^{6},\phi^{4}$ to the attractor point $n_{s}=1-{2\over N_{e}}$, $r={12\over N^{2}_{e}}$ is very fast, the attractor point is reached already at $\xi=O(1)$. Meanwhile for the models $\phi^{2},\phi,\phi^{2/3}$ this limit is reached only for relatively large $\xi$. Thus, for these models, it makes sense to consider a more natural possibility $\xi=O(1)$. This consideration will be important for us later, see section 13.1.